OpenStudy (anonymous):

find the general solution: yy'=-x+((x^2)+(y^2))^0.5

OpenStudy (jamesj):

Note first that (y^2)' = 2yy' and (x^2)' = 2x Thus (y^2 + x^2)' = 2yy' + 2x Hence if we let u = x^2 + y^2, we can rewrite your equation as 1/2 . u' = u^(1/2) Now you can solve that equation by separation of variables.